OpenStudy (anonymous):

helpp please? B^2H^5x= T^3K E^4R Y

OpenStudy (anonymous):

question?

OpenStudy (anonymous):

its over divison line i need to solve for x please

OpenStudy (anonymous):

\[\frac{B^2H^5x}{E^4R}=\frac{T^3K}{Y}\] ?????

OpenStudy (anonymous):

if yes \[B^2H^5x=\frac{T^3K*E^4R}{Y}\] \[x=\frac{T^3K*E^4R}{Y}-B^2H^5\]

OpenStudy (anonymous):

sorry the last step is wrong answer is\[x=\frac{T^3K*E^4R}{Y*B^2H^5}\]

OpenStudy (anonymous):

thanks a lot !!